how many milliliters of water must be added to 250 ml of 25% salt solution to reduce its concentration to 10%
Answer by stanbon(75887)

how many milliliters of water must be added to 250 ml of 25% salt solution to reduce its concentration to 10%
------------------
Equation:
water + water = water
x + 0.75*250 = 0.90(x+250)
----
Multiply thru by 100 to get:
100x + 75*250 = 90x + 90*250
10x = 15*250
x = 375 ml (amt. of pure water needed)
====================
Cheers,
Stan H.
====================